For the quadratic equation, find the value of the discriminant and give the number of real solutions.
3x^2+8x=0

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)
Answer by ikleyn(52781)

.
Referring to the general form of a quadratic polynomial/equation


    p(x) = ax^2 + bx + c = 0 


the formula for the discriminant is 


    d = b^2 - 4ac.


In your case,  a = 3, b = 8, c = 0.


Calculate the discriminant on your own.


If d > 0, there are 2 real solution.

If d = 0, there is only one real solution  ("two real solutions merge into one").

If d < 0,  there are no real solution. (There are two complex solutions instead).


Having this instructions from me, you do the rest.

--------------------
On solving quadratic equations see the lessons
    - Introduction into Quadratic Equations
    - PROOF of quadratic formula by completing the square
in this site.

My other lessons on solving quadratic equations in this site are
    - HOW TO solve quadratic equation by completing the square - Learning by examples
    - Solving quadratic equations without quadratic formula
    - Who is who in quadratic equations
    - Using quadratic equations to solve word problems
    - HOW TO solve the problem on quadratic equation mentally and avoid boring calculations
    - Using Vieta's theorem to solve qudratic equations and related problems
    - OVERVIEW of lessons on solving quadratic equations

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Quadratic equations".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.